UT 17-December-2019

First half TSIP (Chan,LUCI) / Second half ARGOS-Com or Contingency Engineering
  • OSA: D.G.Huerta
  • ISp: G. Bechetti
  • Mtn Manager: M. Wagner
  • ISA: D.Thompson, OKuhn (Tucson)
  • AO: S.Walsh (Tucson)
  • Special Guest Stars: G.Rahmer (Tucson)
UT 17-December-2019

TSIP (Chan,LUCI)

See Nightlog UT20191217 on Google docs for info
Problems near end with AGw4 --> LUCI2/AGw4 offline

OSCO

Skipped because of poor seeing

Transform Collection

Skipped because of poor seeing

ARGOS

UT09:15
Problems opening dust shutter on SX/LM2
Waking MWagner to manually open shutter
...Shutter is open

LBTO Eng/D

Data on Arp 143 taken:
  • Working from obs2 /home/lbto/Service_Observing/Eng_Dec2019
  • Arp 143 collimation script was track mode. Olga fixed it from OT file (Coll.xml)
  • 0.73" from guider images & wfs

Heater appears to have blown circuit breaker on S wall of RemoteOps rm507-3 and rm507-4 (6-headed computers) lost power Plugged into another wall directly with extension cord Rebooted rm507-3

Arp143_ARGOS_L1.xml

UT10:54:47 preset sent, runAO failed

UT10:58:56 preset sent, failed ... NGS not found

UT11:02:30 preset sent

0056 1x[26x2.51s] without high order Zs, 5.0 4.2 pixels 0057 sky 330,-330" offset
0058 source, high gains turned on during exp
0059 source
0060 sky
0061 source - partially high gains off, don't use for measurement
0062 source - high gains off, fwhm 4.8 pix
0063 sky
0064 source
0065 sky
0066 source
11:25 paused for satellite closure
11:29 Shane doing another beam sharpening
11:30 Shane notes he is still seeing ripples with high gain
11:33 resuming LUCI1 script
0067 sky
0068 source
0069 source --- during this one vibration, loops opened
0070 sky --- loops still open
0071 source --- loops still open

11:41 when loops are open, the script should stop, but it did not
0072-0074 are junk - still taking data although configuring for transform

11:45 Configuring for L1 transform data collection. Changed LZP and stopped/started PCS.

Stone_E was too low, but Stone_I was OK. Decided on Stone_I_m93.list with FeII and exptime=4 since the on-axis star is bright. But while collimating, the preset was canceled. It looks like the AZ torques were railed. There is a problem with the AZ frame.

No useful data were taken for the L1 transform and AGw4 problems prevented even attempting the L2 transform confirmation.

-- DavidThompson - 17 Dec 2019
Topic revision: r2 - 17 Dec 2019, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback